Calculate the time taken for 40kC to pass through a component when a current of 5.0A goes through it


1
Expert's answer
2021-03-04T11:53:59-0500

"Q=40kC=40\times10^3\space C"

"I=5A"

By definition of electric charge,

"Q=I\times time\\\Rightarrow time=\dfrac{Q}{I}\\\Rightarrow time=\dfrac{40\times10^3}{5}=8\times10^3\space s"